OpenStudy (anonymous):

Show an equation and a solution for the problem. At the coal mine, it takes the older conveyor belt 20 hours to move one day's coal output. The new belt can do it in 15 hours. If the old belt began to move the day's output and the new belt started as well an hour later, how long would it take to move it? Round to the nearest tenth.

OpenStudy (anonymous):

9.143 hrs

OpenStudy (anonymous):

Old belt can complete 1/20 th of work in one hour and new belt can complete 1/15 th of the work in 1 hour thus both can complete \[\frac{1}{20}+\frac{1}{15} = (\frac{7}{60})\] in one hour and as old belt started one hour before so work left = \[1- \frac{1}{20} = \frac{19}{20}\] there fore together the two belts can complete the remaining work in \[\frac{19}{20}\times \frac{60}{7} = \frac{57}{7} hrs\] therefore total time to complete the work = \[1 + \frac{57}{7} = \frac{64}{7} = 9.142 hrs\]

OpenStudy (saifoo.khan):

sheg is helping people these days!

OpenStudy (anonymous):

b4 i dint used to help

OpenStudy (saifoo.khan):

Right.

OpenStudy (anonymous):

fir mein kya karta tha mere bhai????

OpenStudy (saifoo.khan):

Tafreeiyan! ;)

OpenStudy (anonymous):

haahahhaa.........achha hai boss

OpenStudy (anonymous):

by d way mujhe iska matlab nahi pata hai!!!

OpenStudy (saifoo.khan):

LOL. great.